Two charges are on the x axis: A −16μC charge is at the origin and a 4μC charge is at x = 2cm. Find the magnitude and direction of the electric field at the point P, which is 3cm along the +x axis.


1
Expert's answer
2021-04-23T07:28:51-0400

Answer

Electric field due to −16μC charge

"E_1=\frac{9\times10^9\times16\times10^{-6}}{0.03\times0.03}(-i) \\=16\times10^7(-i)N/C"

Electric field due to 4μC charge

"E_1=\frac{9\times10^9\times4\times10^{-6}}{0.01\times0.01}(i) \\=36\times10^7(i)N/C"

So resultant is

"E=E_1+E_2\\=20\times10^7(i)N/C"

In direction +x axis.